OpenStudy (anonymous):

if f(x)=3x-1 and g(x)=4x+3 find and simplify the composition g[f(x)]

OpenStudy (anonymous):

f(x)=3x-1 g(f(x))=4(3x-1)+3 =12x-4+3 =12x-1